OpenStudy (christos):

I know that we can say P(A or B) = P(A) + P(B) - P(A and B) But can we say P(A and B) = P(A) + P(B) - P(A or B) ? also can we say that P(A' and B') = 1 - P(A and B) , P(A' or B') = 1 - P(A or B)

OpenStudy (anonymous):

no because "or" means multiply; "and" means add. Adding and Multiplying are not interchangeable

OpenStudy (christos):

I saw it in a paper

OpenStudy (anonymous):

okay...? what does that have to do with anything

OpenStudy (christos):

idk .. just wonderin

OpenStudy (anonymous):

oh lol. Well there is your answer

OpenStudy (christos):

also can we say that P(A' and B') = 1 - P(A and B) , P(A' or B') = 1 - P(A or B)

OpenStudy (anonymous):

I dont think so... primes (the little ') are very different

OpenStudy (christos):

' doesnt indicate prime , they are the complements

OpenStudy (anonymous):

well in calculus, the ' indicates prime and that was my first thought

OpenStudy (perl):

but you could simplify ( A and B ) ' = A' or B'

OpenStudy (perl):

that is true from set theory

OpenStudy (perl):

$$ \Large {P(A \cup B) = P(A) + P(B) - P(A\&B) \\~\\ \Large P(A \& B ) = P(A) \cdot P(B|A) \\~\\ P(A ' ) = 1 - P(A) \\ \therefore \\ \Large P[(A \& B)' ] = 1 - P( A\&B) \\ \Large P[(A \cup B)' ] = 1 - P( A\cup B) \\ \therefore \\ \large P[(A \& B)' ] = P[ A' \cup B' ] = P(A' ) + P(B') - P( A' \& B' ) \\ \large P[(A \cup B)' ] =P[ A' \& B' ] = P(A' )\cdot P(B' | A ' ) } $$

OpenStudy (perl):

some authors use \( \Large \cap \) for `and`

OpenStudy (perl):

@mastermindkakashi 'or' implies add in the formula. 'and' means multiply

OpenStudy (anonymous):

ah okay thanks for clarifying @perl though you still can't change them up

OpenStudy (perl):

right, they have separate formulas. substitution has to be done consistently

OpenStudy (kainui):

Actually the answer is very simple Start here, each of these things can be thought of as simply representing numbers and what you can do with numbers works perfectly fine with them, so we can do regular algebra on it. P(A or B) = P(A) + P(B) - P(A and B) So we can just add P(A and B) to both sides to get P(A or B) +P(A and B) = P(A) + P(B) Now subtract P(A or B) from both sides and we get the equation you were looking for. P(A and B) = P(A) + P(B) - P(A or B) I think the best way to visualize this particular instance if you really want to is with venn diagrams.

OpenStudy (christos):

P(A' or B') = 1 - P(A or B) ? P(A' and B') = 1- P(A and B) ? @Kainui @perl

OpenStudy (perl):

no christos, that is false (generally)

OpenStudy (perl):

$$ \Large A' ~or~ B' \neq (A ~or~ B ) ' \\ \Large \color{blue}{\text {But...}\\} \\ \Large A' ~or~ B' = (A ~and~ B ) ' \\ \Large A' ~and~ B' = (A ~or~ B ) ' $$ Therefore: $$ \Large P(A' ~or~ B') =P( (A ~and~ B ) ') = 1 - P( A ~and~ B ) \\~\\ \Large P(A' ~and~ B') =P( (A ~or~ B ) ') = 1 - P( A ~or~ B ) $$

OpenStudy (christos):

thank you